I have tried to work this system can you please tell me if it is corect.
4x-12y= 5
-x+ 3y=-1
11y=-4

10x + 2y = 7
5x + 3
This is as far as I know to go with this one please help me. Thank You so much

Answer by stanbon(48545)

4x-12y= 5
Add 12y to both sides to get:
12y=4x-5
Divide both sides by 12 to get:
y = (1/3)x-(5/12)
===============
10x + 2y = 7
Subtract 10x from both sides to get:
2y = -10x+7
Divide both sides by 2 to get
y = -5x + (7/2)
===============
Cheers,
Stan H.